# Survival and invasion dynamics in cell populations: an analytical framework for threshold behaviour in nonlinear age-structured models

### v1

@@ Abstract
Cell populations invade through a combination of proliferation and motility. Proliferation depends on the internal timing of cell division: how long cells take to complete the cell cycle. This timing varies substantially within (and across) cell types, creating age structure where cells at different times since their last division have different propensities to divide. Classical mathematical models of cell spreading treat division as memoryless and predict exponential cell-cycle-time distributions. Lineage tracing, by contrast, reveals peaked, gamma-like distributions that indicate a maturation delay leading to a fertility window. This gap motivates a modelling framework that incorporates age-dependent cell division rates while retaining analytical tractability.
We address this through a moment-hierarchy framework that tracks time since cell division, with age resetting to zero at division. The framework yields explicit formulae for steady-state age distributions, cell-cycle-time distributions, and invasion speeds. For age-independent rates, we recover classical Fisher–KPP. Three fundamental principles emerge. First, age structure systematically reduces a population’s carrying capacity and narrows the viable parameter range for positive steady states. Second, classical linear theory overestimates invasion speeds; the true minimal speed is slower when division is age-dependent. Third, the parameter condition for population survival is identical to the condition for a positive invasion speed.
Relevance to Life Sciences. Our study addresses a fundamental question in wound healing, tissue regeneration, and tumour invasion: how the timing of cell division regulates whether populations can persist and how fast they spread? By deriving explicit formulae for cell-cycle-time distributions, we link single-cell measurements of division timing to population-level outcomes—thresholds in division and death rates that determine population persistence and the rate of invasion. These formulae enable direct parameter estimation from lineage-tracing data, demonstrated by fitting division-time measurements from cells under control and perturbed conditions. The results show that interventions that prolong the cell cycle shift populations toward extinction. The relationship between division timing and invasion speed provides a mechanistic basis for interpreting competitive outcomes in collective cell migration and tissue dynamics.
Mathematical Content. We formulate an age-structured transport-renewal equation for a motile and proliferating cell population, where cell age tracks time since last division and resets to zero upon division. Cell division and death are modelled as distinct processes, each dependent on cell age and population density. We derive a generalised moment system that connects the total population to its age-weighted moments. This reduction preserves the essential structure of the model while providing a finite-dimensional, tractable system. We classify models by their rate functions—constant, exponentially decaying, or gamma-type—representing age-homogeneous, early-biased, and delayed division. For each case, this framework yields explicit expressions for the population survival condition, the steady-state age and cell-cycle-time distributions, and the minimal invasion speed, all derived without closure approximations.
@@ 1 Introduction
We develop an analytical framework for collective cell invasion that links single-cell division timing to population survival and invasion. Models are classified by their rate functions—constant, exponentially decaying, or gamma-like. Integration with respect to age yields a set of integro–differential equations where age structure enters through weighted integrals. This finite integro–differential formulation can be analysed directly, without closure approximations. Division and death are treated as separate processes, each dependent on cell age and local density. We treat non-spatial and spatial dynamics. In the non-spatial setting, we obtain explicit formulae for steady-state age distributions, cell-cycle-time distributions, and the parameter conditions for population survival. In the spatial setting, we derive expressions for the minimal wave speed and recover the classical Fisher–KPP limit when division and death are age-independent. To our knowledge, this is the first deterministic analysis within this class of nonlinear age-structured PDEs to provide explicit cell-cycle-time distributions together with survival conditions and invasion criteria. It addresses a gap previously approached only through stochastic or discrete-time models [yates_2017, belluccini_2022]. We show that the shape of the division kernel controls both population survival and invasion speed. Biologically, the results account for the gamma-shaped division-time distributions observed in lineage-tracing data and demonstrate how longer division times slow invasion. The framework links measurable quantities to the underlying division and death mechanisms.
@@ 2.1 Integro–differential reduction
where $B(x,t)=u(0,x,t)$ is given by the renewal condition (3). When division and death rates depend only on the total population, the system closes in terms of $P(x,t)$ alone; see Lemma 2.1 (proof in Appendix A). This corresponds to logistic-type dynamics (cf. Case 1).
@@ 5 Travelling waves and invasion thresholds
We now examine travelling-wave solutions of the spatial ($\kappa>0$) models to assess how age structure modifies invasion speed. Classical Fisher–KPP theory predicts a single wave speed determined by the net growth rate. In the context of age-structured models, two distinct linear predictions emerge. The first, $c_{\mathrm{lin}}$, arises from the moment-system approximation and coincides with the Fisher–KPP prediction. The second, $c_{\min}$, obtained from the age-structured formulation, is systematically smaller. This discrepancy arises because the moment-system linearisation assumes uniform decay across all ages at the wave front, whereas the age-structured model preserves differential decay where younger cells dominate the front, and their lower division potential slows invasion. Moreover, the condition for a positive wave speed exactly matches the non-spatial survival threshold, linking persistence and invasion within a unified framework.
@@ 5.2 Age-structured analysis and minimal wave speeds
The minimal wave speed $c_{\min}$ is the infimum of $c>0$ over $\lambda>0$ satisfying (42). For tractable rate forms, (42) yields $c_{\min}$ analytically. Otherwise, it reduces to a root-finding problem that is straightforward to compute numerically.
@@ 5.3 Minimal wave speeds across model cases
In summary, the upper bound $c_{\mathrm{lin}}$ arises from linearising the moment system at the extinction state, assuming uniform decay across all age classes. The exact minimal speed $c_{\min}$ accounts for the age-structured profile at the leading edge and satisfies $c_{\min}\leq c_{\mathrm{lin}}$, with equality only when death and division rates are age-independent. We call invasion conditions the parameter conditions derived by requiring $c_{\mathrm{lin}}>0$, which coincide with the non-spatial survival thresholds for all cases (Table 2 (non-spatial) and Table 5 (spatial)). When invasion conditions hold, travelling waves exist with positive minimal speed; when they fail, populations cannot invade regardless of initial spatial configuration. And that is because population dies out.
@@ 6 Discussion
Appendix A Proof of Lemma 2.1
Proof A.1.
Integrate (2) with respect to $a$ with $u(\infty,x,t)=0$ to give
Appendix B Mathematical and biological background
A fundamental constraint in cell proliferation is the need to complete a sequence of preparatory phases before division can occur. This introduces a delay between a cell’s birth (via division of a mother cell) and its ability to divide again. Experimentally observed division-time distributions are typically peaked and right-skewed, rather than exponential, reflecting this maturation delay. Classical models of population dynamics track total cell density and assume a constant division rate independent of time since the last division. This assumption yields exponentially distributed division times and has enabled major progress in understanding population-level phenomena but does not capture the time delay constraint. To incorporate this biological constraint, we introduce cell age—the time since the last division—as a measurable proxy for cell-cycle progression. The framework then tracks how the population’s age distribution shapes collective outcomes such as survival and spatial spread.
The dynamics are formalised by a PDE describing the cell density $u(a,x,t)$ as a function of age $a$, spatial position $x$, and time $t$. The model incorporates four biological processes: random cell motion; cell ageing at unit rate (a newborn cell at time $t$ reaches age $a$ at time $t+a$); death at rate $\mu(a,P)$, which may depend on both age (e.g., senescence) and local crowding $P$; and division at rate $\beta(a,P)$, which typically increases with age (reflecting maturation) and decreases with crowding (reflecting contact inhibition). Each division event produces two daughter cells of age zero. This age-structured population equation provides a minimal framework linking single-cell dynamics to population-level behaviour. The key idea is that reproduction depends on the distribution of times since the last division, introducing a memory effect: the current division rate depends on the entire age structure, not only on total population size.
A central analytical challenge is the model’s circular dependency: newborn cells (age zero) arise only from older cells dividing, so the division rate at any moment depends on the current age structure. But the age structure itself depends on past division rates. When division and death rates depend on both age structure and density, this creates a feedback loop between population history and current growth. We address this by deriving a reduced system based on weighted integrals of the age distribution, including the total population density $P(x,t)$ and related moments that summarise the population structure. This reduction preserves the essential influence of cell age and yields tractable equations. From these, we obtain explicit, testable expressions for population survival conditions, steady-state age distributions (the fraction of cells at each age in a population at steady-state), and cell-cycle-time distributions (the distribution of division times in a population at steady-state). These results enable parameter estimation directly from lineage-tracing data without requiring full numerical simulation of the PDE system.
We also analyse travelling-wave solutions that describe expanding fronts of cells invading unpopulated regions. A key question is how rapidly the population spreads into free space. Classical theory with exponential division times predicts a wave speed $c_{\mathrm{lin}}=2\sqrt{\kappa r}$, where $\kappa$ is the random motility coefficient and $r$ the net growth rate. In our age-structured framework, the actual speed $c_{\mathrm{min}}$ satisfies $c_{\mathrm{min}}\leq c_{\mathrm{lin}}$. The biological interpretation is that the cell cycle imposes a cost on spatial spread. Cells at the front of the wave are young and thus unlikely to divide immediately. Because local proliferation drives population invasion, this transient loss of reproductive potential reduces the growth rate at the wave front. As a result, we observe a slower invasion speed than predicted by memoryless models. This finding implies that classical, memoryless models systematically overestimate invasion rates when division potential is not homogenoeus in a cell population.
Appendix C Further reading
Cell-cycle regulation and division timing
  • [1] D. O. Morgan, The Cell Cycle: Principles of Control, New Science Press in association with Oxford University Press, London, 2007.
[…]
  • [3] V. K. Gupta and O. Chaudhuri, Mechanical regulation of cell-cycle progression and division, Trends in Cell Biology, 32 (2022), pp. 773–785.
  • [4] J. J. Tyson and B. Novák, Models in biology: lessons from modeling regulation of the eukaryotic cell cycle, BMC Biology, 13 (2015), art. 46.
Age-Structured Population Models and Spatial Dynamics
  • [1] M. Iannelli and F. Milner, The Basic Approach to Age-Structured Population Dynamics, Lecture Notes on Mathematical Modelling in the Life Sciences, Springer, Dordrecht, 2017.
  • [2] H. R. Thieme, Mathematics in Population Biology, Princeton Series in Theoretical and Computational Biology, Princeton University Press, Princeton, NJ, 2003.
  • [3] J. A. J. Metz and O. Diekmann (eds.), The Dynamics of Physiologically Structured Populations, Lecture Notes in Biomathematics, vol. 68, Springer-Verlag, Berlin, 1986.
  • [4] J. D. Murray, Mathematical Biology. I: An Introduction, 3rd ed., Interdisciplinary Applied Mathematics, vol. 17, Springer-Verlag, New York, 2002.
  • [5] J. D. Murray, Mathematical Biology. II: Spatial Models and Biomedical Applications, 3rd ed., Interdisciplinary Applied Mathematics, vol. 18, Springer-Verlag, New York, 2003.
  • [6] O. Diekmann, M. Gyllenberg, and J. A. J. Metz, Steady-state analysis of structured population models, Theoretical Population Biology, 63 (2003), pp. 309–338.

### v2

@@ Abstract
Cell populations invade through a combination of proliferation and motility. Proliferation depends on the internal timing of cell division: how long cells take to complete the cell cycle. This timing varies substantially within (and across) cell types, creating age structure where cells at different times since their last division have different propensities to divide. Classical mathematical models of cell spreading treat division as memoryless and predict exponential cell-cycle-time distributions. Lineage tracing, by contrast, reveals peaked, gamma-like distributions that indicate a maturation delay leading to a fertility window. This gap motivates a modelling framework that incorporates age-dependent cell division rates while retaining analytical tractability. We address this through a moment-hierarchy framework that tracks time since cell division, with age resetting to zero at division. The framework yields explicit formulae for steady-state age distributions, cell-cycle-time distributions, and invasion speeds. For age-independent rates, we recover classical Fisher–KPP. Three fundamental principles emerge. First, age structure systematically reduces a population’s carrying capacity and narrows the viable parameter range for positive steady states. Second, classical linear theory overestimates invasion speeds; the true minimal speed is slower when division is age-dependent. Third, the parameter condition for population survival is identical to the condition for a positive invasion speed.
@@ 1 Introduction
We develop an analytical framework for collective cell invasion that links single-cell division timing to population persistence and invasion. Models are classified by their rate functions—constant, exponentially decaying, or gamma-like. Integration with respect to age yields a set of integro–differential equations where age structure enters through weighted integrals. This finite integro–differential formulation can be analysed directly, without closure approximations. Division and death are treated as separate processes, each dependent on cell age and local density. We treat non-spatial and spatial dynamics. In the non-spatial setting, we obtain explicit formulae for steady-state age distributions, cell-cycle-time distributions, and the parameter conditions for population persistence. In the spatial setting, we derive expressions for the minimal wave speed and recover the classical Fisher–KPP limit when division and death are age-independent. To our knowledge, this is the first deterministic analysis within this class of nonlinear age-structured PDEs to provide explicit cell-cycle-time distributions together with survival conditions and invasion criteria. It addresses a gap previously approached only through stochastic or discrete-time models [yates_2017, belluccini_2022]. We show that the shape of the division kernel controls both population persistence and invasion speed. Biologically, the results account for the gamma-shaped division-time distributions observed in lineage-tracing data and demonstrate how longer division times slow invasion. The framework links measurable quantities to the underlying division and death mechanisms.
@@ 2.1 Integro–differential reduction
where $B(x,t)=u(0,x,t)$ is given by the renewal condition (3). When division and death rates depend only on the total population, the system closes in terms of $P(x,t)$ alone; see Lemma 2.1. This corresponds to logistic-type dynamics (cf. Case 1).
@@
Proof 2.2.
Integrate (2) with respect to $a$ with $u(\infty,x,t)=0$ to give
@@ 5 Travelling waves and invasion thresholds
We now examine travelling-wave solutions of the spatial ($\kappa>0$) models to assess how age structure modifies invasion speed. Classical Fisher–KPP theory predicts a single wave speed determined by the net growth rate. In the context of age-structured models, two distinct linear predictions emerge. The first, $c_{\mathrm{lin}}$, arises from the moment-system approximation and coincides with the Fisher–KPP prediction. The second, $c_{\min}$, obtained from the age-structured formulation, is systematically smaller. This discrepancy arises because the moment-system linearisation assumes uniform decay across all ages at the wave front, whereas the age-structured model preserves differential decay where younger cells dominate the front, and their lower division potential slows invasion. Moreover, the condition for a positive wave speed exactly matches the non-spatial survival condition, linking persistence and invasion within a unified framework.
@@ 5.2 Age-structured analysis and minimal wave speeds
The minimal wave speed $c_{\min}$ is the smallest $c>0$ for which there exists $\lambda>0$ satisfying (43). For tractable rate forms, (43) yields $c_{\min}$ analytically. Otherwise, it reduces to a root-finding problem that is straightforward to compute numerically.
@@ 6 Discussion
See pages - of supplement.pdf